Two charges lie on positive x-axis. Charge A (3.0 x 10^-9 C) is 3.0 cm from the origin and Charge B is 5.0 cm from the origin. (-4.0 x 10^-9 C). What is the total force exerted by these two charges on Charge C (6.0 x 10^-9 C) located at the origin?


1
Expert's answer
2022-03-28T14:24:58-0400

Force


"F_1=\frac{kq_1q_3}{r^2}=\frac{9\times10^9\times3\times10^{-9}\times6\times10^{-9}}{0.03^2}=1.8\times10^{-4}N"

Force


"F_1=\frac{kq_3q_2}{r^2}=-\frac{9\times10^9\times4\times10^{-9}\times6\times10^{-9}}{0.05^2}=0.864\times10^{-4}N"

"F_{net}=F_1-F_2"


"F_{net}=(1.8-0.864)\times10^{-4}N=0.936\times10^{-4}N"